From GlueXWiki
Jump to: navigation, search

May 22, 2014 FDC+CDC meeting


  1. Gas system status
  2. DAQ (Beni)
  3. CDC testing (Mike)
  4. Engineering (Bill)
  5. Electronics (Chris, Nick)
  6. FDC alignment [1],[2](Lubomir)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Naomi (CMU), Eugene, Beni, Dave, Chris, Nick, Simon, Mike, and Lubomir (JLab).

Gas system

- The gas system is ready and, on Thursday morning, we started purging step by step the elements in the gas room. For the weekend we will be flushing also the piping down to the Hall bypassing the detectors.

- It turned out there's Neoprene material in the regulators. Narciso will contact the company to see if one can replace the Neoprene parts with Viton. In principle the area of gas contact is very small and most likely will not cause outgassing problems.


- Beni: the modules needed for the trigger (except the TDC to record scintillator timing) are installed in the trigger crate and all the optical connections are done. There's a problem with the fast readout of the fADC125. Both Beni and Sasha could not read them. It turned out Beni was using slow readout so far in the mini-DAQ system.

CDC status

- Beginning of the week Beni and Mike started flushing the CDC with 40/60 pre-mix bottle at 1 l/min. HV of one sector at 2150V. Mike sees the first signals, though with low amplitudes, assumes due to oxygen contamination.

- Mike: mainly problems with the HV modules. With the help of Nick trying to replace some bad modules to have full system working. After replacing a module without turning off the power and then restarting the HV crate the hardware limits changed by up to few hundred vols!


- Bill (not at the meeting) is looking for a camera to attach to the CMM to scan the cathode at JLab. Simple web camera doesn't work, it also has to have remote focusing.

- Fluorinert level went down by ~2 cm in one month. Dave and Casey were trying the find the leakage with a sniffer without success. Casey will make a tube to sniff inside the magnet. We are going to weigh the quantity the is needed too refill the chiller and estimate the cost. Bill also proposed to put the chiller on a scale.


- After Chris checked all the channels in the Beni's, Fernando concluded that only three of them have problems in the detector (that we knew already about) and further studies will be done with the fADC125 pulser when the firmware will available.

FDC alignment

- Lubomir: first iteration of the alignment of the third package using cosmics done. The corrections to the rotations and pitch values, and the offsets of the cathodes are shown in a table linked above. They correspond to about 0.5mm uncertainties when positioning the frame on the cathodes which is reasonable. As Eugene requested, Lubomir will estimate the errors of these corrections.